### v4

# CBAM-EfficientNetV2 for Histopathology Image Classification using Transfer Learning and Dual Attention Mechanisms

Abstract
Breast cancer histopathology image classification is critical for early detection and improved patient outcomes. This study introduces a novel approach leveraging EfficientNetV2 models, to improve feature extraction and focus on relevant tissue regions. The proposed models were evaluated on the BreakHis dataset across multiple magnification scales (40X, 100X, 200X, and 400X). Among them, the EfficientNetV2-XL with CBAM achieved outstanding performance, reaching a peak accuracy of 98.96% and an F1-score of 98.31% at 400X magnification, outperforming state-of-the-art methods. By integrating Contrast Limited Adaptive Histogram Equalization (CLAHE) for preprocessing and optimizing computational efficiency, this method demonstrates its suitability for real-time clinical deployment. The results underscore the potential of attention-enhanced scalable architectures in advancing diagnostic precision for breast cancer detection.
Index Terms:
Transfer Learning, Modality Specific, EfficientNet, CBAM, Histopathology, Image Classification, Deep Learning, Breast Cancer.
I Introduction
Automating histopathological analysis has become a vital research area in medical imaging due to its potential to significantly reduce the diagnostic workload of pathologists and enhance the consistency and reproducibility of results. Histopathology—the microscopic study of tissue abnormalities—is fundamental in diagnosing diseases such as cancer, but manual analysis is labor-intensive, subjective, and prone to variability between observers. As a result, computer-aided diagnosis (CAD) systems have gained traction, with the aim of augmenting pathologists’ abilities to efficiently and accurately classify tissue samples [1]. Despite substantial progress, CAD systems continue to face significant challenges, particularly in handling the variability inherent in histopathological images. Differences in magnification levels, staining protocols (e.g., Hematoxylin and Eosin, or H&E staining), and tissue morphology introduce variability that complicates the generalization of standard deep learning models [2].
Traditional convolutional neural networks (CNNs), while successful in many computer vision tasks, often struggle to generalize across these variations in histopathological data. Specifically, CNNs are designed with fixed receptive fields and are limited in their ability to capture fine-grained details that vary across multiple scales [3]. The heterogeneity of breast cancer histopathology images—marked by different magnification levels and complex cellular structures—poses unique challenges that demand more sophisticated architectures. To address this, researchers have turned to hybrid models that integrate attention mechanisms with CNNs, which have demonstrated superior performance in medical imaging tasks by enhancing feature extraction.
Attention mechanisms such as the Convolutional Block Attention Module (CBAM) have been adapted to improve image classification by enabling the network to focus dynamically on the most informative regions of an image. In histopathology, these mechanisms are particularly effective in highlighting regions critical to cancer diagnosis, such as irregular nuclei shapes, abnormal cellular arrangements, and other key tissue structures that may indicate malignancy. By applying attention, the network can dynamically recalibrate feature maps along spatial and channel dimensions, improving its ability to capture both local and global image features [4]. This is especially valuable in distinguishing between benign and malignant tissue, where small but significant differences in morphology must be detected.
EfficientNet [8], a state-of-the-art CNN architecture known for its compound scaling strategy, is particularly well-suited for medical image classification tasks due to its ability to balance model depth, width, and resolution. EfficientNet achieves superior performance with fewer parameters, making it computationally efficient without sacrificing accuracy. By integrating attention mechanisms with EfficientNet, we aim to build models that can more effectively classify breast cancer histopathology images while remaining scalable to real-world clinical applications. This hybrid approach addresses the need for models that can operate across varying image resolutions and magnification levels, ensuring robustness and high accuracy in diverse settings.
II Literature Review
Hybrid models have gained significant attention in breast cancer classification due to their ability to combine the strengths of multiple architectures. A multi-model assembling strategy that integrates local and global branches, reducing redundancy and enhancing feature representation. This approach has achieved high accuracy on standard breast cancer datasets, demonstrating the potential of hybrid architectures. Similarly, Togacar et al. [5] developed BreastNet, a model that employs attention mechanisms along with hypercolumn techniques to improve the precision of identifying malignant regions. The inclusion of attention modules has been shown to substantially improve performance, especially in complex tasks like cancer detection.
The use of attention mechanisms in medical imaging has proven highly effective. Woo et al. [4] introduced CBAM, which applies sequential attention along both the channel and spatial dimensions. This method allows networks to highlight critical regions in complex medical images, contributing to substantial performance improvements in classification tasks involving histopathology images. Beyond CBAM, Self-Attention [6] has been adapted from its initial use in natural language processing to capture contextual dependencies across different regions of an image. This has proven particularly useful in medical image analysis, where understanding the relationship between local and global image features is essential for accurate classification.
Recent advancements in hybrid attention mechanisms have significantly improved the capability of deep learning architectures to focus on salient features in complex datasets. For instance, works like Transformer-CNN hybrids [6] have shown remarkable success in combining global context capture through self-attention with local feature extraction inherent in CNNs, making them highly effective for tasks such as lesion segmentation and tissue classification in histopathology.
Another notable development is the Multi-Scale Attention Networks (MSANs), which leverage hierarchical attention mechanisms to capture fine-grained details and high-level contextual information simultaneously. These architectures have proven particularly effective for analyzing multi-resolution data, as demonstrated by Zhang et al. [35], who employed MSANs for high-resolution cellular analysis in breast cancer histopathology, achieving state-of-the-art performance.
Furthermore, Deformable Attention Mechanisms have recently gained traction for their adaptability in handling heterogeneous medical datasets. He et al. [36] introduced deformable attention layers that dynamically adjust receptive fields, outperforming traditional attention mechanisms in scenarios with high inter-patient variability, such as BreakHis and CAMELYON17 datasets.
Hybrid approaches like these emphasize the potential of combining localized feature emphasis (via channel/spatial attention) with global dependency modeling, bridging a critical gap in medical imaging where both local abnormalities and global patterns must be identified. Our approach, integrating CBAM with EfficientNetV2, aligns with this trend but distinguishes itself by optimizing computational efficiency and scalability for clinical deployment, as evidenced by its performance on the BreakHis dataset.
III Methodology
This section outlines the detailed methodology employed in this study for classifying breast cancer histopathology images using Hybrid EfficientNet models integrated with advanced attention mechanisms. Our focus is on enhancing model performance through a robust preprocessing pipeline, modality-specific transfer learning, and a well-defined training strategy.
III-A Model Architecture
Our proposed model is based on EfficientNetV2-XL which use compound scaling to optimize accuracy and efficiency, making them ideal for handling the variability in histopathological images, such as differences in cellular structures and staining quality.
To enhance feature extraction, we integrate the Convolutional Block Attention Module (CBAM). CBAM combines channel and spatial attention, allowing the model to focus on informative features and key regions within the image. This is particularly beneficial in histopathology, where precise identification of cancerous structures is crucial. The EfficientNet backbone extracts initial features, and CBAM refines these representations before passing through fully connected layers for classification.
This integrated architecture, illustrated in Fig. 1, optimizes both efficiency and accuracy in breast cancer histopathology classification.
III-B CBAM Architecture
In this section, we detail the proposed modality-specific CBAM-EfficientNet architecture designed for the classification of H and E stained breast histopathology images. This architecture enhances the traditional EfficientNet framework by integrating the Convolutional Block Attention Module (CBAM), allowing for improved focus on relevant features within the input images.
Figure 1: CBAM
III-B1 Channel Attention Module
The Channel Attention Module (CAM) is pivotal in our architecture for identifying significant features across the channel dimension. This module utilizes the interdependencies among channels, emphasizing crucial information while suppressing less relevant features. The channel attention process follows these steps:
1. Pooling Operations: The input feature map $F$ is first subjected to average pooling and maximum pooling to capture the global contextual information:
$F_{avg}=\text{AvgPool}(F)\quad\text{and}\quad F_{max}=\text{MaxPool}(F)$
2. Feature Aggregation: The pooled feature maps are then fed into a multi-layer perceptron (MLP) to generate channel attention scores, $M_{c}(F)$, which are defined as:
$M_{c}(F)=\sigma\left(\text{MLP}(F_{avg})+\text{MLP}(F_{max})\right)$
where $\sigma$ denotes the sigmoid activation function.
3. Weighting: The resulting channel attention scores are used to weight the original feature map $F$ to obtain the refined feature representation $F^{\prime}_{c}$:
$F^{\prime}_{c}=M_{c}(F)\odot F$
Here, $\odot$ represents the element-wise multiplication, effectively scaling the original features based on their importance.
Figure 2: Model Architecture
III-B2 Spatial Attention Module
Following the channel attention mechanism, the Spatial Attention Module (SAM) refines the feature maps by focusing on the spatial locations of the features. This module enhances the model’s capability to discern where important features are located in the input image.
1. Feature Map Aggregation: The spatial attention mechanism aggregates the features across the channel dimension by applying both average and maximum pooling, resulting in two feature maps:
$F_{avg\_c}=\text{AvgPool}(F^{\prime})\quad\text{and}\quad F_{max\_c}=\text{MaxPool}(F^{\prime})$
2. Spatial Attention Calculation: These aggregated maps are then concatenated and passed through a convolutional layer to produce the spatial attention map $M_{s}(F^{\prime})$:
$M_{s}(F^{\prime})=\sigma\left(f_{7\times 7}([F_{avg\_c};F_{max\_c}])\right)$
where $f_{7\times 7}$ denotes a convolution operation with a $7\times 7$ kernel.
3. **Feature Refinement**: The refined feature map $F^{\prime}_{s}$ is obtained by applying the spatial attention map to the already refined channel-wise features:
$F^{\prime}_{s}=M_{s}(F^{\prime})\odot F^{\prime}$
Figure 3: Spatial and Channel Attention
III-B3 EfficientNet Backbone
The backbone of our model utilizes the EfficientNet architecture, known for its efficiency in parameter utilization and computational performance. Each EfficientNet variant is constructed using a combination of depthwise separable convolutions and squeeze-and-excitation blocks, allowing for enhanced feature extraction while maintaining low computational costs.
We integrate the CBAM modules at various points in the EfficientNet architecture to enhance the focus on the most relevant features. The output from the final layer is then passed through a Global Average Pooling (GAP) layer, which reduces the dimensionality while preserving important spatial information:
$F_{GAP}=\text{GAP}(F^{\prime}_{s})$
Finally, the output from the GAP layer is fed into fully connected layers for classification tasks, ensuring that the model captures both local and global features effectively.
This combined approach of CBAM and EfficientNet leverages both channel and spatial attention mechanisms, providing a robust framework for the classification of breast histopathology images while maintaining computational efficiency
III-C Datasets
For our experiments, we utilized the BreakHis dataset, which consists of 9,109 breast histopathology images sourced from 82 patients. The dataset contains images at four magnification levels: 40X, 100X, 200X, and 400X, categorized into benign and malignant classes. To enhance model generalizability, we also employed additional datasets such as ICIAR 2018 and the PCam dataset for pre-training. The ICIAR 2018 dataset includes a variety of histopathology images labeled for benign and malignant conditions, while the PCam dataset provides a substantial number of whole-slide images with binary labels indicating metastatic tissues.
III-D Data Preprocessing
To address the challenges posed by histopathology images, we implemented a comprehensive hybrid preprocessing pipeline. The following steps were undertaken:
  • Zero Padding: We applied zero padding to handle edge pixels effectively, ensuring that convolutional operations do not lose critical information at the borders of images. This technique adds additional rows and columns around the image, extending beyond its boundary.
  • Median Filtering: To reduce additive noise while preserving the structural integrity of the tissue, we employed a median filter. This non-linear filter replaces each pixel value with the median of its neighborhood values, effectively removing noise without blurring edges, which are vital for accurate feature extraction.
  • Contrast Limited Adaptive Histogram Equalization (CLAHE): Given the uneven staining in histopathology images, CLAHE was utilized to enhance local contrast and improve weak boundary detection. By dividing the image into small tiles and applying histogram equalization to each, we improve the visibility of critical features within the images.
  • Normalization: Image pixel values were normalized to the range [0, 1] to enhance model convergence during training. This ensures uniformity in the input data distribution, which is crucial for the performance of deep learning models.
  • Data Augmentation: To mitigate the challenges of limited sample sizes, extensive data augmentation was performed on the BreakHis dataset. Techniques such as rotation, flipping, zooming, and brightness adjustment were employed to artificially increase the dataset size and introduce variability, thereby reducing the risk of overfitting.
Figure 4: Data Augmentation and preprocessing techniques
III-E Training and Hyperparameter Tuning
The training of the models was conducted using PyTorch on CUDA, enabling efficient computation. We utilized an initial learning rate of 0.001 and employed Adam optimizers to determine which provided better convergence and generalization.
Key hyperparameters were carefully tuned:
  • Dense Layers: Experiments were conducted with two and three dense layers to identify the optimal configuration for the final classifier.
  • Activation Functions:ReLU activation functions was used at the last classification layer.
  • Kernel Initializers: We experimented with different kernel initializers, including the normal initializer and He initialization, to determine their effects on the convergence rate during training.
  • Dropout Regularization: Dropout was incorporated in the fully connected layers to prevent overfitting, particularly given the limited availability of labeled data in medical imaging tasks with a rate of 0.4.
  • Batch Size and Epochs: We tested various batch sizes (16, 32, and 64) and finalized with 32 and trained the models for a maximum of 100 epochs, applying early stopping based on validation loss to prevent overfitting.
Given the inherent limitations of training deep learning models with small datasets, we employed a modality-specific transfer learning approach. This strategy leverages pre-trained models on large, similar datasets to enhance feature extraction capabilities for our specific task.
We fine-tuned the EfficientNet models, initially pre-trained on diverse cancer datasets (PCam, ICIAR 2018, and others) rather than the ImageNet dataset, to reduce the domain gap inherent in medical imaging tasks. By tailoring the transfer learning process to domain-specific data, we aimed to improve model robustness and classification accuracy on the BreakHis dataset.
TABLE I: Performance of Base Models Across Magnification Scales
Model Scale Accuracy (%) F1 Score (%)
EfficientNetV2-XL (Base) 40X 86.78 84.23
100X 87.89 85.34
200X 88.23 86.67
400X 89.12 87.56
EfficientNetV2-L (Base) 40X 85.67 83.13
100X 86.89 84.23
200X 87.13 85.56
400X 87.89 86.23
III-F Evaluation Metrics
The performance of our proposed models was assessed using a suite of evaluation metrics: accuracy, precision, recall, and F1-score. These metrics are crucial in medical applications where both false negatives and false positives can have severe implications for patient care. Moreover, the performance was evaluated across different magnification levels (40X, 100X, 200X, and 400X) to analyze how well the models generalize to various scales of tissue detail.
Figure 5: Visualization of Attention Mechanisms in Histopathological Images
The experiments were conducted using PyTorch as the deep learning framework backend. The models were trained and tested on an NVIDIA RTX 4080 GPU, providing substantial computational power for the deep learning tasks involved. The system was equipped with 16GB of RAM to efficiently handle the large dataset and facilitate faster processing.
Hyperparameters are critical variables that govern the training process of convolutional neural networks (CNNs). While the networks learn the relationships between inputs and outputs, optimizing hyperparameters is essential to enhance model performance. We performed extensive experiments to identify the optimal hyperparameters for our final training.
We evaluated two different optimizers: Adam (Adaptive Moment Estimation), with learning rate ( $0.001$. The SoftMax classifier was employed for all classification experiments, providing probability scores for each class label.
Binary cross-entropy was used as the loss function due to the binary nature of our classification problem, represented mathematically as follows:
[…]
The fully connected layer incorporated a ReLU activation function and consisted of 256 hidden neurons, followed by a dropout layer with a probability of $0.4$ to mitigate overfitting.
Figure 6: ROC-AUC Curve of EfficientNetV2-XL with CBAM
IV Experimental Results and Discussions
IV-A Results
We employed a modality-specific transfer learning strategy for the classification task and compared it with various state-of-the-art CNN architectures, as shown in Table II.
Initially, we conducted experiments on two different models without transfer learning, as presented in Table 3. Subsequently, we evaluated state-of-the-art models pre-trained on ImageNet, as summarized. The performance of our proposed modality-specific strategy, incorporating transfer learning.
As shown in Table 2, our proposed EfficientNetV2-XL + CBAM significantly outperformed various established models, including DenseNet201 and Deep CNNs. Notably, at higher magnification levels, this model achieved an outstanding accuracy of 98.96% and an F1-score of 98.31% at 400X magnification, highlighting its ability to accurately classify breast cancer histopathology images
The EfficientNetV2-XL + CBAM demonstrated substantial performance improvements over the VGG16 + VGG19 combination across all magnifications. While the VGG-based models delivered competitive accuracy, their computational requirements were significantly higher. For instance, the VGG16 + VGG19 models required a Tesla K80 GPU with 24GB of VRAM to accommodate their large parameter count and high computational demand.
In contrast, the EfficientNetV2-XL + CBAM was trained on a more modest NVIDIA RTX 4080 GPU with just 16GB of VRAM. Despite the lower memory availability, the EfficientNetV2-XL architecture demonstrated exceptional computational efficiency, leveraging depthwise separable convolutions, squeeze-and-excitation (SE) blocks, and the Convolutional Block Attention Module (CBAM).
TABLE II: Comparison of Accuracy Across Models and Magnification Factors
Method Model MG Factor Accuracy (%)
Erfankhan Hamed, et al LBP 40X 88.30
100X 88.30
200X 87.10
400X 83.40
Daniel Lichtblau, et al AlexNet 40X 81.61
100X 84.47
200X 86.67
400X 83.15
Yun Gu, et al DCMM 40X 95.62
100X 95.03
200X 97.04
400X 96.31
Nahid, et al Deep CNN 40X 90.00
100X 91.00
200X 91.00
400X 90.00
Togacar Mesut, et al Deep CNN 40X 97.99
100X 97.84
200X 98.51
400X 95.88
Yan Hao, et al DenseNet 201 40X 96.75
100X 95.21
200X 96.57
400X 93.15
Pin Wang, et al FE-BkCapsNet 40X 92.71
100X 94.52
200X 94.03
400X 93.54
VGG16 + VGG19 Hameed et al 40X 94.44
100X 97.61
200X 98.70
400X 98.96
Proposed Hybrid 2024 40X 97.11
(EfficientNetV2-XL + CBAM) 100X 98.04
200X 98.25
400X 98.96
V Conclusion
In this study, we propose a Hybrid EfficientNetV2 models integrated with CBAM, Self-Attention, and Deformable Attention mechanisms for breast cancer histopathology classification. The EfficientNetV2-XL model, with approximately 208 million parameters and the inclusion of the CBAM module, offers a more efficient and accurate solution compared to previous state-of-the-art models like VGG16 + VGG19, which have around 282 million parameters. This reduction in parameters enables faster training and reduced memory consumption, allowing the model to be trained on an RTX 4060 GPU with 8GB of memory.
The Hybrid EfficientNetV2-XL + CBAM and previous State-of-the-Art model VGG16 + VGG19 model achieved superior accuracy, however EfficientNetv2-XL has F1-Score of 98.31 and Hybrid VGG16 + VG199 has 97.95% ,therefore EfficientNetV2-XL + CBAM outperforms other models in both accuracy and computational efficiency. By leveraging attention mechanisms, the model effectively focuses on critical regions of histopathological images, making it well-suited for real-time clinical diagnostics. These results demonstrate that the proposed model surpasses the performance of previous state-of-the-art models, highlighting its potential for real-time clinical deployment. Future work may adapt the model to other datasets and further expand its clinical applicability.